OpenStudy (anonymous):

test the series for converge or divergence. n^2e^(-n). I'm not sure what method to use here. I tried integral test...but that didnt seem to work.

OpenStudy (anonymous):

use the ratio test:$$\lim_{n\to\infty}\left|\frac{(n+1)^2e^n}{n^2e^{n+1}}\right|=\lim_{n\to\infty}\left(\frac{n+1}{n}\right)^2\frac{e^n}{e^{n+1}}=\frac1e\lim_{n\to\infty}\left(1+\frac1n\right)^2=\frac1e$$

OpenStudy (anonymous):

\(1/e<1\) hence the series converges

OpenStudy (anonymous):

thanks oldrin! Youre the man. How did you know to use the ratio test?

OpenStudy (anonymous):

oh ok...I just learned this stuff today, so Im still working on it, but thanks for your help!

OpenStudy (anonymous):

oops I did the root test incorrectly:$$\lim_{n\to\infty}\sqrt[n]{n^2e^{-n}}=\frac1e\lim_{n\to\infty}n^{2/n}=\frac1e<1$$

OpenStudy (anonymous):

@Requiem no problem; once you get used to them it's a lot easier to know which test seems adequate

OpenStudy (anonymous):

I'd also suggest you try proving that the tests work :-p

OpenStudy (anonymous):

:P one step at a time oldrin!